# Direct optimization of neoclassical ion transport in stellarator reactors

### v1

@@ Abstract
We directly optimize stellarator neoclassical ion transport while holding neoclassical electron transport at a moderate level, creating a scenario favorable for impurity expulsion and retaining good ion confinement. Traditional neoclassical stellarator optimization has focused on minimizing $\epsilon_{\mathrm{eff}}$, the geometric factor that characterizes the amount of radial transport due to particles in the $1/\nu$ regime. Under expected reactor-relevant conditions, core electrons will be in the $1/\nu$ regime and core fuel ions will be in the $\sqrt{\nu}$ regime. Traditional optimizations thus minimize electron transport and rely on the radial electric field ($E_{r}$) that develops to confine the ions. This often results in an inward-pointing $E_{r}$ that drives high-$Z$ impurities into the core, which may be troublesome in future reactors. In our optimizations, we increase the ratio of the thermal transport coefficients $L_{11}^{e}/L_{11}^{i}$, which previous work has shown can create an outward-pointing $E_{r}$. This effect is very beneficial for impurity expulsion. We obtain self-consistent density, temperature, and $E_{r}$ profiles at reactor-relevant conditions for optimized equilibria. These equilibria are expected to enjoy significantly improved impurity transport properties. We conclude by providing several directions of future research that may help further improve the presented optimization algorithm.
@@ 1 Introduction
Stellarator reactors are expected to achieve relatively high temperatures in their plasma cores, resulting in the electrons and fuel ions exhibiting neoclassical transport characteristics of the long mean free path (lmfp) regime. This regime may be divided into several pieces that can be visualized in figure 12 of reference [1]. The $1/\nu$ regime corresponds to zero normalized $\mathbf{E}_{r}\times\mathbf{B}$ drift velocity, $v_{E}^{*}\equiv\left|E_{r}\right|/vB_{0}=0$, where $\mathbf{E}_{r}=E_{r}(r)\nabla r$ is the radial electric field, $r$ is a flux label, $v$ is particle speed, and $B_{0}$ is a reference magnetic field strength. The radial particle and heat transport in the $1/\nu$ regime can be characterized by the magnitude of the geometric factor $\epsilon_{\mathrm{eff}}$ [2], which has been commonly used as a minimization target for neoclassical stellarator optimizations. If $v_{E}^{*}>0$, two other lmfp regimes exist: the $\nu$ regime at extremely low collisionalities and the $\sqrt{\nu}$ regime at more moderate collisionalities. The thermal transport coefficients for particle and heat fluxes are approximately proportional to $\left|E_{r}\right|^{-2}$ in the former case and $\left|E_{r}\right|^{-3/2}$ in the latter [3]. Roughly speaking, $\mathbf{E}_{r}\times\mathbf{B}$ rotation helps limit the radial excursions of particles from flux surfaces when particle behavior is not entirely dominated by collisional effects. Since $\epsilon_{\mathrm{eff}}$ characterizes transport levels when $v_{E}^{*}=0$, lowering $\epsilon_{\mathrm{eff}}$ lowers the “ceiling” on radial transport for populations in all lmfp regimes and thus reduces overall radial transport. However, because $v_{E}^{*}\approx 0$ only for electrons under expected reactor-relevant conditions, minimizing $\epsilon_{\mathrm{eff}}$ directly minimizes only radial electron transport. One may think that radial ion transport would be larger than radial electron transport — the larger ion masses lead to larger radial excursion of the orbits and therefore larger diffusion coefficients — but the charge separation immediately establishes the aforementioned $\mathbf{E}_{r}$ to make the transport ambipolar [4]. The electric field will point radially inward in this case ($E_{r}<0$). Because the ambipolar $\mathbf{E}_{r}$ is determined by the condition that the net radial current is zero and the ions are the primary contributor to the radial current in (reasonably) low-$\epsilon_{\mathrm{eff}}$ configurations, the $E_{r}<0$ solution of the ambipolarity condition is typically referred to as the “ion root” ambipolar electric field.
@@
In design studies, stellarator reactors have commonly targeted $\epsilon_{\mathrm{eff}}$ values in the range $0.1\%$$1\%$ [5, 6], which is sufficiently small to prevent neoclassical energy losses from being an obstacle to ignition but not small enough to preclude other design objectives from being taken into account. Such reactors are expected to have strong ion root electric fields, so impurities (e.g., partially-ionized tungsten atoms from the divertor and other plasma-facing components) would be driven into the core and cause substantial radiative losses that would likely make continuous reactor operation difficult. The presence of a neoclassical impurity pinch has long been recognized in the literature [7, 8, 9, 10, 11], and several remedies have been proposed. For instance, the temperature gradient of the fuel ions may help counteract the pinch associated with a small ion root electric field [12, 13], but this will likely not be effective if a strong ion root is present throughout the plasma. Furthermore, tangential electric fields may make this effect relatively unhelpful as they often enhance the impurity pinch caused by the radial electric field, especially for high-Z impurities [14, 15, 16, 17]. Buller et al. attempted to directly optimize the tangential electric field (through the electric potential) to prevent inward impurity flux but were only able to reduce it [18].
@@
Another possible solution exists for the impurity transport problem: if $E_{r}>0$ in some region of the plasma, this will provide a thermodynamic force that drives impurities (especially those with high $Z$) radially out of that region. The case $E_{r}>0$ is referred to as the “electron root” solution of the ambipolarity condition. Operating stellarators with an electron root was historically considered superior to ion root operation since the larger $\left|E_{r}\right|$ characteristic of the electron root is favorable for confinement [19, 20]. On the other hand, experimental realizations of the electron root typically involve substantial electron heating that causes the electron temperature to be much higher than the fuel ion temperature [9, 21, 22, 23, 24], which contradicts the desire for reactors to have strong collisional coupling such that plasma thermal energy efficiently drives fusion reactions. This has lead to skepticism that stellarator reactors can feasibly operate with a steady-state electron root.
@@
In this work, we seek to show that the electron root may not be as elusive in reactors as previously thought. Under the assumptions that transport processes are radially local and the drift kinetic equation can be linearized about a Maxwellian, the neoclassical particle flux may be written as
@@
with $\iota$ the rotational transform on a given flux surface, $R_{0}$ a reference major radius, $m_{s}$ the mass of species $s$, $K_{s}=\frac{m_{s}v_{s}^{2}/2}{T_{s}}=\left(\frac{v_{s}}{v_{s,\mathrm{thermal}}}\right)^{2}$, $D_{ij}^{*}$ the monoenergetic transport coefficient evaluated for a given $v_{E}^{*}$, $h_{1}=h_{3}=1$, and $h_{2}=K_{s}$ [3]. While it may initially appear that $L_{11}^{e}/L_{11}^{i}\ll 1$ will always hold given the mass scaling in equation 2, it has been previously noted (to help explain the “impurity hole” phenomenon of the Large Helical Device [25, 26]) that sufficiently large $L_{11}^{e}/L_{11}^{i}$ will cause $E_{r}>0$ [12]. Optimized configurations have also recently been discovered by Beidler and coworkers with extremely small $D_{11}^{*}$ at fuel-ion-relevant $v_{E}^{*}$ and moderate values of $\epsilon_{\mathrm{eff}}$ [27], leading to a large $L_{11}^{e}/L_{11}^{i}$ and therefore a strong electron root in the core under reactor-relevant conditions. The theoretical foundations of this phenomenon are explored in more detail in recent work by Helander et al. [28].
@@
The objective of our work is to follow the suggestion of reference [27] and optimize stellarator magnetic configurations to increase $L_{11}^{e}/L_{11}^{i}$. We control $L_{11}^{e}$ through $\epsilon_{\mathrm{eff}}$ and $L_{11}^{i}$ through an array of $D_{11}^{*}$ evaluated at fuel-ion-relevant $v_{E}^{*}$ and $\nu^{*}\equiv R_{0}\nu/\iota v$ (normalized collisionality) values. Monoenergetic transport coefficients should, in principle, provide a more general target than thermal transport coefficients because they are characteristic of a given magnetic field and do not depend on plasma profiles. Our transport optimizations are entirely neoclassical — because turbulence is often the dominant transport mechanism in optimized stellarators such as Wendelstein 7-X [29], the techniques presented here may only be useful when coupled with turbulence optimization techniques. Indeed, lowering the overall level of turbulence in a reactor may well reduce the outward impurity flux provided by the turbulence and necessitate the presence of another impurity removal mechanism, such as an electron root. We note that large $E_{r}^{\prime}$ (and thus large $\mathbf{E}_{r}\times\mathbf{B}$ shear), which is common in electron root scenarios, may help reduce radial turbulent fluxes [30, 31]. Large $E_{r}^{\prime}$, usually in the ion root region near the last closed flux surface, is known to allow access to improved confinement modes of operation somewhat similar to H-modes in tokamaks [9, 32, 33].
2 General Formulation
$\chi^{2}=\sum_{i}\chi_{i}^{2},$ (3)
$\chi_{i}^{2}=\frac{\left(f_{i}-f_{i}^{\mathrm{target}}\right)^{2}}{\sigma_{i}^{2}},$ (4)
where $f_{i}$ is any quantity that is callable by the STELLOPT routines, $f_{i}^{\mathrm{target}}$ is the desired value for $f_{i}$, and $\sigma_{i}$ is a user-specified inverse weight. VMEC [35] is used as the MHD equilibrium solver in all cases. Differential evolution [36] and a Garabedian boundary representation [37] are utilized to find global minima more reliably except where noted in section 3.1.
The optimizations begin from a fixed-boundary configuration based on the W7-X high-mirror configuration [38] whose volume and on-axis magnetic field strength have been scaled to the reactor-relevant values of roughly 1756 $\mathrm{m^{3}}$ and 5 T, respectively [39]. Stellarator symmetry and five field periods are assumed for all optimizations. The initial value of $\beta$ varies between optimizations as detailed in the next section. Zero bootstrap current is assumed during the MHD calculations within the optimization loop. The core feature of the optimizations is the minimization of an array of $D_{11}^{*}$ with fuel-ion-relevant $v_{E}^{*}$ and $\nu^{*}$. The $D_{11}^{*}$ values are calculated by DKES [40, 41], and a $3\times 3$ array consisting of $\left|E_{r}\right|/v=\left\{5.00\times 10^{-4},2.74\times 10^{-3},1.50\times 10^{-2}\right\}~{}\mathrm{T}$ and $\nu/v=\left\{2.50\times 10^{-6},1.58\times 10^{-5},1.00\times 10^{-4}\right\}~{}\mathrm{m^{-1}}$ is used. (These quantities, rather than their normalized counterparts, are specified as targets in STELLOPT.) The exact values of $\left|E_{r}\right|/v$ and $\nu/v$ seem relatively unimportant — the upper and lower values for each quantity are chosen to be approximately reactor-relevant bounds, and the middle values are the logarithmic means of the upper and lower values. Notably, modifications of $B_{0}$, $R_{0}$, and $\iota$ are relatively small in the optimizations, so explicitly targeting $\left|E_{r}\right|/v$ and $\nu/v$ rather than $v_{E}^{*}$ and $\nu^{*}$ is acceptable. Even nearly doubling $R_{0}$ at the beginning of the Configuration 1 optimization (see section 3.1) made no noticeable difference to the output quality. In addition to $D_{11}^{*}$, we target $\epsilon_{\mathrm{eff}}=1\%$ (calculated by NEO [2]) in all optimizations except where noted in section 3.1. All neoclassical objectives are calculated for the $\rho=r/a=\{0.25,0.5,0.75\}$ flux surfaces. MHD objectives such as plasma aspect ratio, elongation, and others are also included in the optimizations. Section 3 gives tabulated objective information for each optimization presented.
Once optimized, the transport characteristics of each configuration are analyzed using DKES and NTSS [42]. NTSS performs flux-surface-averaged, time-dependent transport calculations (including momentum correction techniques [43]) to determine self-consistent profiles for density, temperature, radial electric field, bootstrap current, and other quantities on 51 radial grid points for a given equilibrium. It should be emphasized that the equilibrium itself is not updated during the transport calculations, but NTSS calculates bootstrap current using the $D_{31}^{*}$ database provided by DKES and updates the rotational transform profile by utilizing susceptance matrices [44]. Reactor volumes are rescaled to 1450 or 1900 $\mathrm{m^{3}}$ and on-axis magnetic field strengths are rescaled to 5 or 6 T to achieve an alpha power of 600 MW because this approximately corresponds to a 3 GWth reactor. (We note that the electron root is expected to be more difficult to achieve as the magnetic field strength increases [3, 31], which may need to be taken into account if our methods are used alongside those for turbulence optimization.) Deuterium and tritium density profiles are fixed with on-axis values for each in excess of $10^{19}$ $\mathrm{m^{-3}}$ and core gradients relatively flat in $\rho$ — see figure 4 for details. NTSS uses the D-T cross-section to calculate the birth rate of alpha particles, which for simplicity are assumed to slow down on the flux surface of their birth. This determines the helium density, and quasineutrality then determines electron density. We assume there are no heavy impurities in the plasma. The temperature profiles for each species are initialized with roughly parabolic shapes in $\rho$, and the initial on-axis temperatures vary between analyses as described in the next section. A very simple turbulent transport model is included based on fits to W7-AS data. It takes the form [45]
$Q_{s}^{\mathrm{turb}}=-\chi_{\mathrm{turb}}n_{s}T^{\prime}_{s},$ (5)
$\Gamma_{s}^{\mathrm{turb}}=-D_{\mathrm{turb}}n^{\prime}_{s},$ (6)
where $Q_{s}^{\mathrm{turb}}$ is the radial energy flux of species $s$ due to turbulence. Similarly to reference [27], we set $\chi_{\mathrm{turb}}=0.0065(P/\mathrm{MW})^{3/4}/\left(n_{s}/10^{20}~{}\mathrm{m^{-3}}\right)~{}\mathrm{m^{2}s^{-1}}$ and $D_{\mathrm{turb}}=0.0003(P/\mathrm{MW})^{3/4}/\left(n_{s}/10^{20}~{}\mathrm{m^{-3}}\right)~{}\mathrm{m^{2}s^{-1}}$ where $P=P_{\alpha}-P_{\mathrm{Br}}$, $P_{\alpha}$ is the power deposited by fast ions, and $P_{\mathrm{Br}}$ is the power radiated due to Bremsstrahlung. No external heating is considered. NTSS determines the full profile of the radial electric field by solving a diffusion-like differential equation with hysteretic behavior [46] that minimizes the generalized heat production rate of the plasma [47]. This is necessary because the ambipolarity condition generally produces either one or three possible solutions for $E_{r}$ [19, 48, 49] and a discontinuous $E_{r}$ profile is nonphysical. In the case of three solutions, one is the ion root, one is the electron root, and one is an unstable root that is not physically realized. We choose the “diffusion coefficient” in the diffusion-like equation to be 2 $\mathrm{m^{2}/s}$ such that the transition region from the electron root to the ion root is roughly 5 cm wide, which is reasonable in light of global neoclassical simulations [31]. Specifying a detailed startup scenario is beyond our scope, so we initialize NTSS runs by instructing the software to “select” the electron root solution if it exists. This situation is relevant for reactors that primarily utilize electron heating for startup. NTSS steps forward in time until it converges on a steady-state, self-consistent solution for the plasma profiles. We focus only on making the electron root realizable from a neoclassical standpoint and ensuring the optimized configurations can support the electron root during steady-state operation. (Configurations with insufficiently optimized ion confinement, such as scaled W7-X high-mirror, may be artificially seeded with an electron root during high-density operation but will transition to an ion root relatively quickly.)
3 Results
Figure 1: Toroidal cross-sections of the last closed flux surface equally spaced in the toroidal angle across one half-field-period for (1(a)) Configuration 1 (section 3.1), (1(b)) Configuration 2 (section 3.2), (1(c)) Configuration 3 (section 3.3), and (1(d)) scaled W7-X high-mirror.
Figure 2: Scans of $D_{11}^{*}$ versus $\nu^{*}$ with various $v_{E}^{*}$ at $\rho\approx 0.35$ for (2(a)) Configuration 1 (section 3.1), (2(b)) Configuration 2 (section 3.2), (2(c)) Configuration 3 (section 3.3), and (2(d)) W7-X high-mirror. The upper and lower variational bounds on the $D_{11}^{*}$ values are calculated by DKES and bracketed by the “$+$” symbols, but these “error bars” are too small to see in most cases. The dotted lines indicate $D_{11}^{*}$ for a tokamak with the same aspect ratio, rotational transform, and (average) elongation as the given configuration for the same $\rho$. Note that the light blue and purple curves, roughly corresponding to fuel-ion-relevant $v_{E}^{*}$, are far more depressed in the optimized configurations (2(a)2(c)) than in W7-X high-mirror. The characteristic gap between the (fuel-ion-relevant) high-$v_{E}^{*}$ curves and (electron-relevant) $v_{E}^{*}=0$ curves is also much larger in the optimized configurations than in W7-X high-mirror.
Figure 3: The geometric factor $\epsilon_{\mathrm{eff}}$ for the configurations outlined in section 3 and W7-X high-mirror. These calculations are performed on the optimized equilibria without taking effects from the NTSS modeling (such as modifications to bootstrap current, $\beta$, etc.) into account. The large values of $\epsilon_{\mathrm{eff}}$ in Configuration 1 as $\rho\rightarrow 1$ are likely numerical error — the high elongation of this configuration means many Boozer modes are required to fully describe the outermost surfaces, leading any codes that rely on Boozer coordinates to have questionable results there.
Figure 4: NTSS results for the steady-state density and temperature profiles of Configuration 1 (4(a)+4(d)), Configuration 2 (4(b)+4(e)), and Configuration 3 (4(c)+4(f)). The profiles of radial electric field (4(g)), $L_{11}^{e}/L_{11}^{i}$ (4(h)), bootstrap current (4(i)), and $\iota$ (4(j)) are also included. Note that the results for Configuration 1 are suspect for $\rho\gtrsim 0.8$ — see the caption of figure 3 or the text in section 3.1 for an explanation.
Table 1: Plasma parameters of the configurations described in section 3. The “Initial $P_{\alpha}$” and “Initial $P_{Br}$” entries refer to the alpha power and bremsstrahlung power, respectively, at the beginning of the NTSS simulation for the given configuration. The mirror ratio does not change throughout a simulation as the field strength is a user-specified parameter. All other entries refer to the end of the relevant NTSS simulation. $\tau_{E}$ is the energy confinement time, and $\tau_{E}^{\mathrm{ISS04}}$ is calculated using the ISS04 scaling expression [50].
Quantity Configuration 1 Configuration 2 Configuration 3
Mirror Ratio (%) 42.1 27.8 26.9
$\beta\right>_{\mathrm{Vol}}$ (%) 4.94 4.43 3.14
Initial $P_{\alpha}$ (MW) 482.6 777.0 495.6
Final $P_{\alpha}$ (MW) 489.2 628.9 601.5
Initial $P_{Br}$ (MW) 43.8 38.3 33.5
Final $P_{Br}$ (MW) 72.7 73.7 72.7
$\tau_{E}$ (s) $2.13=1.30\tau_{E}^{\mathrm{ISS04}}$ $2.25=1.40\tau_{E}^{\mathrm{ISS04}}$ $2.34=1.23\tau_{E}^{\mathrm{ISS04}}$
3.1 Configuration 1
Table 2: Optimization objectives used to produce Configuration 1 (section 3.1). Note that stages 1 and 2 are performed using STELLOPT, while SciPy optimization routines coupled with VMEC and SFINCS are used for stage 3. Note also that the $L_{ij}^{s}$ values listed in stage 3 refer to the direct inputs for SFINCS, so they are unitless and have a different sign convention than the $L_{ij}^{s}$ in equations 1 and 2. All targets except the aspect ratio are evaluated on the $\rho=\{0.25,0.5,0.75\}$ flux surfaces.

### v2

@@ Abstract
We directly optimize stellarator neoclassical ion transport while holding neoclassical electron transport at a moderate level, creating a scenario favorable for impurity expulsion and retaining good ion confinement. Traditional neoclassical stellarator optimization has focused on minimizing $\epsilon_{\mathrm{eff}}$, the geometric factor that characterizes the amount of radial transport due to particles in the $1/\nu$ regime. Under expected reactor-relevant conditions, core electrons will be in the $1/\nu$ regime and core fuel ions will be in the $\sqrt{\nu}$ regime. Traditional optimizations thus minimize electron transport and rely on the radial electric field ($E_{r}$) that develops to confine the ions. This often results in an inward-pointing $E_{r}$ that drives high-$Z$ impurities into the core, which may be troublesome in future reactors. In this work, we increase the ratio of the thermal transport coefficients $L_{11}^{e}/L_{11}^{i}$, which previous research has shown can create an outward-pointing $E_{r}$. This effect is very beneficial for impurity expulsion. We obtain self-consistent density, temperature, and $E_{r}$ profiles at reactor-relevant conditions for an optimized equilibrium. This equilibrium is expected to enjoy significantly improved impurity transport properties.
@@ 1 Introduction
Traditional neoclassical stellarator optimizations have focused on minimizing the geometric factor $\epsilon_{\mathrm{eff}}$, which characterizes the particle and heat transport of the $1/\nu$ regime [1]. The $1/\nu$ regime is a subset of the long mean free path (lmfp) regime with zero $\mathbf{E}_{r}\times\mathbf{B}$ drift velocity, $v_{E}^{*}\equiv\left|E_{r}\right|/vB_{0}=0$, where $\mathbf{E}_{r}=E_{r}(r)\nabla r$ is the radial electric field, $r$ is a flux label, $v$ is particle speed, and $B_{0}$ is a reference magnetic field strength. Stellarator reactors are expected to achieve temperatures in excess of $10~{}\mathrm{keV}$ in their plasma cores, resulting in the electrons and fuel ions exhibiting neoclassical transport behavior characteristic of the lmfp regime. Because the mass of the electrons is much lower than that of the ions, their characteristic thermal velocity is much higher. Thus we may reasonably approximate $v_{E}^{*}\approx 0$ for electrons in a reactor, indicating that their neoclassical transport behavior will be very near the “ideal” $1/\nu$ regime behavior characterized by $\epsilon_{\mathrm{eff}}$. For the heavier fuel ions, on the other hand, $v_{E}^{*}\approx 0$ is a poor approximation, indicating that they will not be near the $1/\nu$ regime and $\epsilon_{\mathrm{eff}}$ is not expected to accurately characterize their neoclassical transport behavior. Instead, we must consider two additional lmfp regimes which are visualized in figure 12 of reference [2]: the $\nu$ regime at extremely low collisionalities and the $\sqrt{\nu}$ regime at more moderate collisionalities. Particle and heat fluxes are approximately proportional to $\left|E_{r}\right|^{-2}$ for the former and $\left|E_{r}\right|^{-3/2}$ for the latter [3]. Roughly speaking, $\mathbf{E}_{r}\times\mathbf{B}$ rotation helps limit the radial excursions of particles from flux surfaces when particle transport behavior is not entirely dominated by collisional effects. Since $\epsilon_{\mathrm{eff}}$ characterizes transport levels when $v_{E}^{*}=0$, lowering $\epsilon_{\mathrm{eff}}$ lowers the “ceiling” on radial transport for populations in all lmfp regimes and thus reduces overall radial neoclassical transport. Reduced electron transport is the source of the overall transport reduction in low-$\epsilon_{\mathrm{eff}}$ configurations, so an inward-pointing $\mathbf{E}_{r}$ ($E_{r}<0$) develops to prevent charge separation [4], or equivalently, to make the net radial current zero (the so-called “ambipolarity condition”). Intuitively speaking, the radial electric field serves to “hold in” the fuel ions. The case when $E_{r}<0$ is known as the “ion root” ambipolar electric field.
@@
In design studies, stellarator reactors have commonly targeted $\epsilon_{\mathrm{eff}}$ values in the range $0.1\%$$1\%$ [5, 6], which is sufficiently small to prevent neoclassical energy losses from being an obstacle to ignition but not small enough to preclude other design objectives from being taken into account during the optimization process. Such reactors are expected to have strong ion root electric fields. This provides a thermodynamic force that drives impurities (e.g., partially-ionized tungsten atoms from the divertor and other plasma-facing components) toward the core. If techniques to minimize the outward turbulent electron and fuel ion transport in stellarators (see, for example, references [7, 8]) continue to advance, devices designed with these techniques will presumably have significantly reduced turbulent impurity transport as well. At some point, the “pinch” effect of the ion root electric field may be able to overpower the “screening” effect of turbulence. In this case, impurities would accumulate in the core and cause substantial radiative losses that would likely make continuous reactor operation difficult. The presence of a neoclassical impurity pinch has long been recognized in the literature [9, 10, 11, 12, 13], and several remedies have been proposed. For instance, the temperature gradient of the fuel ions may help counteract the pinch associated with a small ion root electric field [14, 15], but this will likely not be effective if a strong ion root is present throughout the plasma. Furthermore, tangential electric fields may make this effect relatively unhelpful as they often enhance the impurity pinch caused by the radial electric field, especially for high-Z impurities [16, 17, 18, 19]. Buller et al. attempted to directly optimize the tangential electric field (through the electric potential) to prevent inward impurity flux but were only able to reduce it [20].
@@
Another possible solution exists for the impurity transport problem: if $E_{r}>0$ in some region of the plasma, this will provide a thermodynamic force that drives impurities (especially those with high $Z$) radially out of that region. When an outward-pointing radial electric field solves the ambipolarity condition, it is referred to as the “electron root” ambipolar electric field. Operating stellarators with an electron root was historically considered superior to ion root operation since the larger $\left|E_{r}\right|$ characteristic of the electron root is favorable for confinement [21, 22]. On the other hand, experimental realizations of the electron root typically involve substantial electron heating that causes the electron temperature to be much higher than the fuel ion temperature [11, 23, 24, 25, 26] and/or hollow density profiles [11, 23, 26]. These situations are unlikely to be favorable or realizable in reactor scenarios due to the presence of turbulence, even if it is reduced through optimization. Furthermore, both situations contradict the desire for reactors to have strong collisional coupling (particularly in the core) such that plasma thermal energy efficiently drives fusion reactions. This has lead to skepticism that stellarator reactors can feasibly operate with a steady-state electron root.
@@
In this work, we seek to show that a steady-state electron root may be realistic in stellarator reactors. Under the assumptions that transport processes are radially local and the drift kinetic equation can be linearized about a Maxwellian, the neoclassical particle flux may be written as
@@
with $\iota$ the rotational transform on a given flux surface, $R_{0}$ a reference major radius, $m_{s}$ the mass of species $s$, $K_{s}=\frac{m_{s}v_{s}^{2}/2}{T_{s}}=\left(\frac{v_{s}}{v_{s,\mathrm{thermal}}}\right)^{2}$, $D_{ij}^{*}$ the monoenergetic transport coefficient evaluated for a given $v_{E}^{*}$, $h_{1}=h_{3}=1$, and $h_{2}=K_{s}$ [3]. While it may initially appear that $L_{11}^{e}/L_{11}^{i}\ll 1$ will always hold given the mass scaling in equation 2, it has been previously noted (to help explain the “impurity hole” phenomenon of the Large Helical Device [27, 28]) that sufficiently large $L_{11}^{e}/L_{11}^{i}$ will cause $E_{r}>0$ [14]. Intuitively, this is because large $L_{11}^{e}/L_{11}^{i}$ signifies that radial electron transport is greater (in some sense) than radial ion transport, necessitating an outward-pointing radial electric field to maintain ambipolarity. Mathematically, raising $L_{11}^{e}/L_{11}^{i}$ increases the value of $E_{r}$ necessary to maintain ambipolarity when the density and temperature gradients are negative — see equation (3) of reference [14]. Optimized configurations have also recently been discovered by Beidler and coworkers with moderate values of $\epsilon_{\mathrm{eff}}$ and extremely small $D_{11}^{*}$ at fuel-ion-relevant $v_{E}^{*}$ [29], leading to a large $L_{11}^{e}/L_{11}^{i}$ and therefore a strong electron root in the core under reactor-relevant conditions. The theoretical foundations of this phenomenon are explored in more detail in recent work by Helander et al. [30].
@@
The objective of our work is to follow the suggestion of reference [29] and optimize a stellarator magnetic configuration to increase $L_{11}^{e}/L_{11}^{i}$. We control $L_{11}^{e}$ through $\epsilon_{\mathrm{eff}}$ and $L_{11}^{i}$ through an array of $D_{11}^{*}$ evaluated at fuel-ion-relevant $v_{E}^{*}$ and $\nu^{*}\equiv R_{0}\nu/\iota v$ (normalized collisionality) values. Rather than driving $\epsilon_{\mathrm{eff}}$ as low as possible and relying on the ion root that develops to confine the fuel ions, we directly minimize the fuel-ion-relevant $D_{11}^{*}$ such that the fuel ions (rather than the electrons) become the rate-determining species. This encourages the development of an electron root to “hold in” the electrons, rather than an ion root to “hold in” the fuel ions. Monoenergetic transport coefficients should, in principle, provide a more general target than thermal transport coefficients because they are characteristic of a given magnetic field and do not depend on plasma profiles. Our transport optimization is entirely neoclassical — because turbulence is often the dominant transport mechanism in optimized stellarators such as Wendelstein 7-X [31], the techniques presented here may only be useful when coupled with turbulence optimization techniques. We note that large $E_{r}^{\prime}$ (and thus large $\mathbf{E}_{r}\times\mathbf{B}$ shear), which is common in electron root scenarios, may help reduce radial turbulent fluxes [32, 33]. Large $E_{r}^{\prime}$, usually in the ion root region near the last closed flux surface, is known to allow access to improved confinement modes of operation somewhat similar to H-modes in tokamaks [11, 34, 35].
2 Formulation
Table 1: Optimization objectives used to produce the example configuration. The neoclassical targets are evaluated on the $\rho=\{0.25,0.5,0.75\}$ flux surfaces. The rotational transform is evaluated on the $s=\rho^{2}=\{0,0.5,1\}$ flux surfaces.